OpenStudy (anonymous):

Easy integral question: What is the integral from pi to 2pi: 0 dx?

OpenStudy (anonymous):

I think it is zero, but I would like some affirmation.\[\int\limits_{\pi}^{2\pi}0 dx\]

OpenStudy (anonymous):

I found the solution thanks to wolfram: http://prntscr.com/3pw9yo